Session: Nov 2004 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Poppi at Mon Nov 8 00:16:48 2004) Success :
no known problems.

Torun (K. Borkowski at Mon Nov 8 06:41:09 2004) Minor failures :
BBC No 3 had wrong attenuation from 18:00 till 18:22 UTC, then replaced with another unit. No other known problems.

Hartebeesthoek (Jonathan Quick at Mon Nov 8 06:59:30 2004) Success :
No known problems

Effelsberg (A. Kraus at Mon Nov 8 08:25:44 2004) Success :

Onsala (Michael Lindqvist at Mon Nov 8 08:57:27 2004) Success :
Problems with the noisecal (i.e, bad tsys-values) during some scans. It did not fire. Can be seen if one compare the tpi/tpical-values where the tsys-valus are bad. Otherwise no known problems.

Westerbork (Tony Foley at Mon Nov 8 10:53:30 2004) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. No known problems. For single dish the operator had to change LO manually

Cambridge (Alastair Gunn at Mon Nov 8 11:11:11 2004) Success :
No known problems.

(Other) (Alastair Gunn at Mon Nov 8 11:11:59 2004) Minor failures :
Used Darnhall telescope. LO was incorrectly set for lower part of the band until about 23:10 UT. Upper part of band was correct. Expect signal in only two VCs. No other known problems.

Effelsberg (ZP, for Dave Graham at Mon Nov 8 15:43:14 2004) Minor failures :
Probably for the first part of n04m1 at Effelsberg the frequency was not right. After 18:49 should be OK.

Noto (P.Cassaro at Wed Nov 10 09:33:15 2004) Success :